# Multidata Causal Discovery for Statistical Hurricane Intensity Forecasting

### v1

@@ 2.3 SHIPS Developmental Dataset
To evaluate how causally informed predictors improve operational forecasting, we extract SHIPS predictors for the same 247 North Atlantic TC cases (2000-2021) used in the ERA5 experiments. Importantly, we only used predictor values from the 00-hour initialization time in the SHIPS developmental dataset. These values are derived from GFS analysis fields, which represent the initial-time atmospheric conditions (i.e., observations assimilated into the GFS model) at forecast hour zero. Unlike GFS forecast fields that estimate future atmospheric states, analysis fields reflect the best available estimate of the atmosphere at the initialization time. This setup places us in a purely statistical—rather than statistical-dynamical—framework, enabling a direct analogy between the SHIPS developmental dataset and the ERA5-derived dataset used for causal predictor discovery. We compare the generalization skill of statistical models trained on the original 21 SHIPS predictors (“original SHIPS”) with models trained on the same predictors augmented with causally selected predictors (“SHIPS+”). These additional predictors were identified using ERA5-based causal discovery experiments but are computed from GFS analysis fields in exactly the same way as the original SHIPS predictors. This setup tests whether the predictors selected in the ERA5 replication setting can improve forecast skill when integrated into the operational SHIPS framework.
@@ 3.1 Causal Feature Selection
We methodologically innovate by adopting the open-source TiGraMITe causal discovery package runge2018causal ; runge2019detecting to find potential predictors of hurricane intensity change at different lead times. TiGraMITe has been widely used to discover causal relationships in climate data kretschmer2016using ; Kretschmer2018 ; samarasingheetal2021 . The multidata PC (M-PC) algorithm, a modified version of the first step of the Peter Spirtes and Clark Glymour Momentary Conditional Independence (PCMCI+) algorithm runge2019detecting ; runge2020discovering , is used to discover time-lagged causal relationships in time-series data. Compared to the more advanced PCMCI+, which aims to learn the full causal graph including contemporaneous causal links, this work uses only the PC algorithm to eliminate potential spurious associations in lagged statistical prediction models iglesias2024causally ; runge2015optimal . The “M” in M-PC indicates the multidata approach, where time series from multiple hurricanes are analyzed together to identify a consistent set of causal predictors. Clear non-causal drivers within the 215-variable ERA5 predictor dataset are detected and excluded through conditional independence tests, resulting in a single set of predictors that remain relevant across all training TCs saranya2023selecting . In our setup, this is done without considering time-lagged variables — each predictor is evaluated at a single time-step for the given forecast lead time. Here we approximate conditional independence relations through a linear partial correlation test. By filtering out non-causal predictors in the statistical models, we aim to improve the generalizability of such models for unseen TC cases.
The M-PC algorithm’s hyperparameters include the minimum and maximum time lags $\left(\tau_{\mathrm{min}},\tau_{\mathrm{\max}}\right)$ and the statistical significance threshold for the partial correlation conditional independence tests $pc_{\alpha}$ to decide on variable removal. We set the minimum and maximum time lag in this study to $\tau_{\mathrm{min}}=4$ steps (1 day) and $\tau_{\mathrm{max}}=20$ steps (5 days) to discover potential novel TC predictors. The M-PC algorithm ranks the strength of the causal relationship between different predictors and TC intensity change by the absolute partial correlation value. Predictors are considered non-causal if the partial correlation between the variable and TC intensity change, conditional on the remaining variables, is non-significant. The names of causal predictors and their relevant time lags are stored in Python dictionaries.
@@ 3.2 Validation procedure
As causal discovery algorithms are sensitive to hyperparameter choice lawrence2021data , the output of the linear causal feature selection framework should not be taken as truth for a complex physical problem such as TC intensity, but rather as a plausible guess of the true causal relationships. Careful cross-validation of the causal predictor lists is necessary to establish their robustness and trustworthiness. We postulate that causal predictors can be validated by the performance of the trained models on the validation data unseen during training. Our working assumption, supported by previous work saranya2023selecting , is that model generalization skill improves when adding causally relevant predictors, and degrades as spurious or redundant features are added. As such, our framework should identify a unique feature set that maximizes validation skill and highest likelihood of being causally related to TC intensification. We use 7-fold cross-validation with 31 storms per split, resulting in $31\times 6=186$ storms for training and 31 storms for validation per fold. This yields multiple candidates of plausible causal relationships in the dataset. For one cross-validation fold, we train a series of regression models based on PC output (causal variable names and relevant time lags) at different statistical significance thresholds ($pc_{\alpha}$). Smaller $pc_{\alpha}$ values typically yield fewer predictors, though not always, which can result in different sets with the same cardinality. We assess the generalization capabilities of the trained MLR models with the coefficient of determination ($R^{2}$) and Pearson’s Correlation Coefficient (PCC).
Figure 1: Multidata causal feature selection methodology. Step 1: Preprocessed spatiotemporal fields for all TC cases form an ensemble of aligned time series, which may contain spurious or non-causal relationships due to autocorrelation or confounding. Step 2: These multivariate time series (training set) are input to the multidata causal discovery algorithm (M-PC), which selects candidate predictors while controlling set size via hyperparameters. Each candidate set is evaluated using cross-validated regression. Step 3: Predictors appearing in at four out of the seven folds are pooled to form the final feature set. The goal is to estimate the portion of the true causal graph that helps predict TC intensity changes.
@@ 3.3 Filtering new potential predictors
Figure 1 shows our framework to discover new predictors using causal discovery algorithms. For all MLR models trained, we calculate the coefficient of determination $R^{2}$ for training, validation, and test sets. For each fold, we choose the model with the highest validation performance ($R^{2}$ and PCC), resulting in a set of 7 best models for analysis (Upper-right panel in Figure 1). Summarizing the $R^{2}$ of all trained models allows us to identify the most generalizable MLR model as the MLR model with the best $R^{2}$ score on the validation TCs. Adding more predictors beyond the set of inputs in the most generalizable MLR model will lead to model overfitting on the training TCs. These variables make the MLR model less generalizable because they constitute spurious associations arising from autocorrelation, indirect effects, or common drivers. To ensure robustness across cases, the final causal variable list presented in subsequent sections is produced by aggregating the variables used by the 7 best models. Only predictors that appear in more than 3 out of 7 best models and are not part of the existing SHIPS developmental predictor list will be shortlisted as new candidate predictors (lower right panel in Figure 1).
@@ 3.4 Experimental Design
Different experiments are conducted to demonstrate the potential applications of causal feature selection algorithms in complementing or enhancing existing statistical models for TC intensity change.
@@ 3.4.1 ERA5-based experiments for predictor discovery
The first set of experiments involves discovering candidate causal predictors from ERA5 reanalysis data that can be incorporated into the operational SHIPS. We designed two experiments to see whether causal feature selection framework can discover such variables and the sensitivity of discovered variables to the PC algorithm’s assumptions. The first experiment (withASSUMPS) assumes that we will not remove any existing SHIPS predictors but only add new ones, representing a least-change case that is more feasible operationally. The second experiment (noASSUMPS) is a free-run, “kitchen-sink” scenario where the algorithm is allowed to overlook existing SHIPS predictors if they do not have strong causal relationships to TC intensity. We switch between the two scenarios in our framework by changing the predefined causal links in the dictionary input fed to a parameter (link_\_assumptions) in the Tigramite package. The link_\_assumptions input in the withASSUMPS experiments is edited to have preexisting causal links between the operational SHIPS predictors and the rate of TC intensity change at all times, regardless of model settings. In contrast, the noASSUMPS experiments do not contain predefined causal links and perform M-PC calculation fairly on all available predictors. Through this experiment, we investigate improving the generalizability of TC statistical models by removing spurious associations in the original SHIPS predictors.
@@ 3.4.2 Validation in an operation-like setting
The second set of experiments assesses the practical value of the causal predictors identified from the ERA5 reanalysis by testing them within the operational SHIPS framework. Specifically, an additional set of six diagnostic variables is extracted from the GFS fields, guided by the most robust predictors shortlisted from the ERA5 discovery experiments. To validate whether these new predictors improve generalization skill in an operational context, we replicate the same experimental setup described previously — comparing models trained with the original SHIPS predictors against models that incorporate the newly suggested predictors. This comparison directly quantifies whether enriching the SHIPS developmental dataset with causally selected predictors yields measurable improvements in forecast performance on unseen TC cases.
@@ 3.5.1 Mapping and Optimization Objective
Parameters are estimated by least squares for consistency with SHIPS and with our linear, partial-correlation–based causal discovery. This yields a deterministic forecast without an explicit predictive distribution. We adopt this deterministic objective because (i) our focus is short lead times where mean-squared error is a standard operational target, (ii) the causal discovery framework deployed here is tailored to mean relationships rather than full conditional distributions, and (iii) maintaining compatibility with SHIPS facilitates a clean comparison in the baseline experiments that follow. Probabilistic forecasting at medium-range and subseasonal horizons is important but out of scope here.
@@ 3.5.2 Baseline Regression models and Baseline Feature Selection Methods
We compare the performance of PC-based causal feature selection models to different baselines to establish the validity of causal discovery methods to suggest useful predictors. Feature selection based on feature correlation and random forest feature importance are used as the chosen linear and nonlinear baselines to compare against the causal feature selection outputs. The correlation baseline is obtained by ranking the correlation between the input variables and rate of TC intensity change. The variable ranking is then used to train linear regression models with increasing complexity in a sequential order. We follow the same process for the random-forest-based baseline, the main difference is that the variable ranking is now the Gini impurity-based feature importance of trained random forest regression models that predict =TC intensity change rates. We repeat these feature selection baselines for the withASSUMPS and noASSUMPS versions of the ERA5 and SHIPS-based experiments.
@@ 3.5.3 Regression architecture
All predictor sets are evaluated using simple multiple linear regression (MLR) models to maintain interpretability and isolate the effect of feature selection. For each experiment, we apply consistent 7-fold cross-validation settings and forecast lead times (24–120 hr), reporting performance on out-of-sample test sets to assess generalizability. We further evaluate the predictor sets using multilayer perceptron (MLP) models, which we select based on their ability to capture nonlinear behaviors in the data that cannot be expressed with MLR models. Each of the MLP models in our study includes a total of 5 layers, wherein each layer includes 512 units (also referred to as neurons). The first three layers rely on a Rectified Linear Unit (ReLU) activation function, the fourth layer instead uses hyperbolic tangent (tanh) to allow the model to output positive and negative intensity changes, and the final layer linearly combines the output of the 4th layer. All of our models were trained to minimize the mean square error (MSE) using the Adam optimizer Kingma2014 with default parameters, a learning rate of 0.001, and with early stopping set to end training if the validation loss is greater than the average of the last 50 epochs once the model has trained for at least 50 epochs. We note that this simplified architecture was chosen with rules of thumb and has not been subjected to a rigorous hyperparameter search given that the objective of these MLPs is not to achieve maximum performance, but rather to provide a simple nonlinear baseline to serve as a comparison to the MLRs.
@@ 4.1 Effectiveness of multidata causal discovery for hurricane intensity prediction
Figure 2: Example results for the 24-hour intensity change forecast (DELV24) from Fold 3 using the SHIPS+ERA5 predictor set for SHIPS predictors. (a) Coefficient of determination $R^{2}$ on training, validation, and test sets plotted against the number of selected predictors, each point corresponding to a different value of the M-PC causal discovery hyperparameter pc_alpha (bottom scale). The vertical dashed line indicates the configuration with the highest validation $R^{2}$. (b) Variable selection abacus: each dot shows the presence of a predictor across the pc_alpha range. Variables are colored by group (e.g., Original SHIPS predictors, Shear, Humidity), vertical dashed line marks the best validation score, and encircled dots highlight the occurrence of new shortlisted predictors for SHIPS.
We apply the multidata PC algorithm to ERA5 reanalysis data exclusively, avoiding inconsistencies from combining different observational sources. For the target variable—the intensity change over various lead times (24 to 120 hours)—we use observed maximum wind speed from IBTrACS to maintain consistency across experiments. We conduct two parallel sets of experiments for each lead time: (1) With SHIPS link assumptions (withASSUMPS): We encode prior knowledge by imposing direct causal links between operational SHIPS predictors and the intensity change target in the TiGraMITe M-PC algorithm. withASSUMPS respects the physical relationships of the original SHIPS model while allowing discovery of additional causal links among other variables. (2) Without SHIPS link assumptions (noASSUMPS or “kitchen-sink”): We allow the algorithm to freely explore causal links among all available ERA5 predictors without any predefined constraints. Figure 2 illustrates an example outcome for the 24-hour intensity change forecast (DELV24) from a representative cross-validation fold (Fold 3) under the kitchen-sink setup. The top panel shows training, validation, and test $R^{2}$ scores as a function of the number of selected predictors, controlled by the M-PC hyperparameter ${pc}_{\alpha}$, which varies from $1.5\ 10^{-4}$ to 0.6. Here, ${pc}_{\alpha}$ represents the significance level used in the conditional independence tests within the PC-stable algorithm. Higher values of ${pc}_{\alpha}$ correspond to a less strict statistical threshold for rejecting independence, allowing more predictors to be included in the causal model whereas a lower ${pc}_{\alpha}$ is more strict, resulting in a more stringent selection of predictors.
@@
The bottom panel presents the variable selection abacus, highlighting predictor presence and groupings across the range of ${pc}_{\alpha}$ values. The encircled dots highlight the predictors included in the final shortlisted set, identified based on their consistent selection frequency across multiple folds. The same procedure was repeated for seven cross-validation folds, all forecast lead times, and both experimental setups. Given the volume of results, Figure 2 presents one representative fold without SHIPS link assumptions, while the best performing folds for lead times of up to 120 hours are provided in the Supplementary Information (Figures S1–S10). Predictors consistently selected across folds and validated configurations form a refined set that complements or improves upon the baseline SHIPS predictors for TC intensity change forecasting.

### v2

@@ 2.3 SHIPS Developmental Dataset
To evaluate how causally informed predictors improve operational forecasting, we extract SHIPS predictors (Table S1) for the same 247 North Atlantic TC cases (2000-2021) used in the ERA5 experiments. We only used predictor values from the 00 hour initialization time in the SHIPS developmental dataset derived from GFS analysis fields, which represent the atmospheric conditions (i.e., observations assimilated into the GFS model) at forecast initialization time. This setup enables direct analogy between the SHIPS developmental dataset and the ERA5-derived dataset used for causal predictor discovery, as both rely on analysis time atmospheric fields within a purely statistical framework. We compare the generalization skill of statistical models trained using the original 21 SHIPS predictors (“original SHIPS”) with models trained on the same predictors augmented with causally selected predictors (“SHIPS+”). These additional predictors are identified through ERA5-based causal discovery experiments but are computed from GFS analysis fields in exactly the same way as the original SHIPS predictors. This setup tests whether the predictors selected in the ERA5 replication setting improve forecast skill when integrated into the operational SHIPS framework.
@@ 3.1 Causal Feature Selection
Causal feature selection is performed using the open-source TiGraMITe causal discovery framework Runge (2018); Runge et al. (2019b) which has been widely used to discover causal relationships from climate data Kretschmer et al. (2016, 2018); Samarasinghe et al. (2021). We employ the multidata PC (M-PC) algorithm, a modified version of the first step of the PCMCI algorithm Runge et al. (2019b); Runge (2020), to discover predictors of TC intensity change while eliminating spurious statistical associations. Unlike PCMCI+, which aims to infer the full causal graph including contemporaneous causal links, this work uses only the PC algorithm for causal feature selection Iglesias-Suarez et al. (2024); Runge et al. (2015). The “M” in M-PC indicates the multidata approach, where an ensemble of time series from multiple TCs is analyzed to identify a single set of causal predictors. Non-causal drivers within the ERA5 predictor set are detected and excluded through conditional independence test using linear partial correlation Ganesh et al. (2023). In our setup, this is done without considering time-lagged variables — each predictor is evaluated at a single time-step for the given forecast lead time. By filtering out non-causal predictors in the statistical models, this approach aims to improve the generalizability of statistical models to unseen TC cases.
The hyperparameters of M-PC include the minimum and maximum time lags $\left(\tau_{\mathrm{min}},\tau_{\mathrm{\max}}\right)$ and the statistical significance threshold for the partial correlation conditional independence tests $pc_{\alpha}$ which determines variable removal. Predictors are ranked according to the absolute value of the partial correlation between each predictor and TC intensity change, conditional on the remaining variables, since the partial correlation test statistic can take both positive and negative values. To assess the robustness of the selected predictors, causal discovery experiments are repeated for each forecast lead time across a wide range of statistical significance thresholds ($pc_{\alpha}\in[10^{-4},\,0.6]$; see SI for the full list) and for multiple single-lag configurations, with $\tau_{\min}=\tau_{\max}\in\{4,8,12,16,20\}$ time steps. For each configuration, predictors exhibiting non-significant conditional dependence are excluded, yielding multiple candidate feature sets across lead times that are subsequently evaluated through cross-validation.
@@ 3.2 Validation procedure
Causal discovery algorithms are sensitive to hyperparameter choice Lawrence et al. (2021), hence, the output of the causal feature selection framework should not be interpreted as truth for a complex physical problem such as TC intensity change. Instead, it may be considered as a plausible approximation of the true causal relationships. Careful cross-validation of the causal predictor lists is necessary to establish their robustness and generalizability. Following previous work Ganesh et al. (2023), we postulate that causally relevant predictors improve model generalization skill, whereas the inclusion of spurious or redundant features degrades performance. By testing the performance of trained models on the validation data unseen during training, our framework should identify a unique feature set that maximizes validation skill and has the highest likelihood of being directly linked to TC intensification. We use storm-based seven-fold cross-validation on the non-test portion of the dataset. Due to year-based partitioning and the removal of storms with no valid samples after preprocessing and alignment, six folds contain 184 training storms and 32 validation storms, while one fold contains 186 training storms and 30 validation storms. Each fold yields multiple candidates of plausible causal relationships, so we train a series of multiple linear regression (MLR) models using predictor sets derived from the PC algorithm at different statistical significance thresholds ($pc_{\alpha}$). Smaller $pc_{\alpha}$ values generally yield fewer predictors, although different thresholds can result in feature sets of same cardinality. We assess the generalization capabilities of the trained MLR models with the coefficient of determination ($R^{2}$).
Figure 1: Multidata causal feature selection methodology. Step 1: Preprocessed spatiotemporal fields for all TC cases form an ensemble of aligned time series, which may contain spurious or non-causal relationships due to autocorrelation or confounding. Step 2: These multivariate time series (training set) are input to the multidata causal discovery algorithm (M-PC), which selects candidate predictors while controlling set size via hyperparameters. Each candidate set is evaluated using cross-validated regression. Step 3: Predictors appearing in minimum four out of the seven folds are pooled to form the final feature set. The goal is to estimate the portion of the true causal graph that helps predict TC intensity changes.
@@ 3.3 Filtering new potential predictors
Figure 1 summarizes the causal feature selection and model evaluation framework. For all trained MLR models, we calculate the $R^{2}$ for training, validation, and test sets. For each cross-validation fold, we choose the model with the highest validation performance ($R^{2}$), resulting in 7 best performing models for analysis (upper-right panel in Figure 1). Comparing validation performance across all trained models allows us to identify the most generalizable MLR configuration. Adding more predictors beyond those chosen by the most generalizable MLR model will lead to model overfitting and make the MLR model less generalizable because they constitute spurious associations arising from autocorrelation, indirect effects, or common drivers. To ensure robustness across cases, the final causal variable list is obtained by aggregating and ranking the variables in the 7 best models. Only predictors that appear in at least four out of seven best models and are not part of the existing SHIPS developmental predictor list will be shortlisted as a candidate predictor (lower right panel in Figure 1).
@@ 3.4 Experimental Design
A hierarchy of experiments is conducted to demonstrate the potential applications of causal feature selection algorithms in complementing or enhancing existing statistical models for TC intensity change. We depict our overall experimental design in Figure 2.
Figure 2: Workflow for causal feature selection, predictor shortlisting, and integration into SHIPS:ERA5-derived variables and ERA5-based replications of SHIPS predictors are used with IBTrACS intensity change as ground truth. Causal discovery is performed across multiple lead times under constrained (withASSUMPS) and unconstrained (noASSUMPS) setups. Predictors shortlisted across lead times based on ranking and performance are integrated into SHIPS+, followed by regression model comparison, SHAP-based interpretation, case-study analysis, and operational SHIPS testing.
@@ 3.4.1 ERA5-based experiments for predictor discovery
The first set of experiments involves identifying candidate causal predictors from ERA5 reanalysis data that can be incorporated into the operational SHIPS. We designed two complementary experiments to assess the algorithms and understand the sensitivity of discovered variables to the assumptions within the causal discovery process. In the first experiment (withASSUMPS), existing SHIPS predictors are retained a priori, and the causal discovery framework is restricted to identifying additional predictors. This represents a least-change and operationally feasible scenario. In contrast, the second experiment (noASSUMPS) is a free-run, “kitchen-sink” that allows the algorithm to freely evaluate all available predictors, including the possibility of excluding existing SHIPS predictors that do not exhibit strong causal relationships with TC intensity change. Through this experiment, we investigate whether removing spurious associations among the original SHIPS predictors improves the generalizability of statistical intensity prediction models. Causal feature selection is however constrained by the properties of the underlying datasets. ERA5 does not fully resolve the inner-core structure of tropical cyclones, and limitations in representing small-scale and azimuthally asymmetric processes are therefore reflected in the candidate predictors, particularly those related to the TC inner core. As a result, some physically relevant variables may not be identified as robust causal predictors within this framework.
@@ 3.4.2 Validation in an operation-like setting
The second set of experiments assesses the practical value of the causal predictors identified using the ERA5 reanalysis by testing them within the operational SHIPS framework. Guided by the most robust predictors identified in the ERA5-based causal discovery experiments, we compute the corresponding predictors from GFS analysis fields and extend the SHIPS developmental dataset including these newly identified causal predictors alongside the original predictors. We then compare models trained using the original SHIPS predictor set with those trained on the augmented predictor set (SHIPS+), using the full SHIPS developmental dataset. This comparison directly quantifies whether enriching the SHIPS developmental dataset with causally selected predictors yields measurable improvements in forecast performance on unseen TC cases.
@@ 3.5.1 Mapping and Optimization Objective
Model parameters are estimated by least squares for consistency with SHIPS and with the linear, partial-correlation-based causal discovery framework. This yields a deterministic forecast without an explicit predictive distribution. We adopt this deterministic objective because (i) our focus is short lead times where mean-squared error is a standard operational target, (ii) the causal discovery framework is tailored to mean relationships rather than full conditional distributions, and (iii) maintaining compatibility with SHIPS facilitates a clean comparison in the baseline experiments.
@@ 3.5.2 Baseline Regression models and Baseline Feature Selection Methods
We compare the performance of PC-based causal feature selection models to different feature selection baselines including lagged correlation and random forest feature importance. The correlation baseline is obtained by ranking the linear correlation between the input variables and rate of TC intensity change. The variable ranking is then used to sequentially train linear regression models with increasing complexity. In random-forest-based baseline, the variable ranking is done with the Gini impurity-based feature importance of trained regression models. We apply these feature selection baselines for the withASSUMPS and noASSUMPS versions across selected lead times for both ERA5 and SHIPS-based experiments.
@@ 3.5.3 Regression architecture
We first evaluate all predictor sets using multiple linear regression (MLR) models to preserve interpretability and isolate the impact of feature selection. For each experiment, we repeat 7-fold cross-validation across forecast lead times (24–120 hr), assessing performance on independent test sets. We further evaluate the predictor sets using multilayer perceptron (MLP) models, which we select based on their ability to capture nonlinear behaviors in the data that are not captured by MLR models. Each of the MLP models in our study includes a total of 5 layers, wherein each layer includes 512 units (also referred to as neurons). The first three layers rely on a Rectified Linear Unit (ReLU) activation function, the fourth layer instead uses hyperbolic tangent (tanh) to allow the model to output positive and negative intensity changes, and the final layer linearly combines the output of the 4th layer. All models were trained to minimize the mean square error (MSE) using the Adam optimizer Kingma and Ba (2014) with default parameters, a learning rate of 0.001, and with early stopping set to end training if the validation loss is greater than the average of the last 50 epochs once the model has trained for at least 50 epochs. This simplified architecture was chosen with rules of thumb and has not been subjected to a rigorous hyperparameter search given that the objective of these MLPs is not to achieve maximum performance, but rather to provide a simple nonlinear baseline for comparison to the MLRs.
@@ 4.1 Effectiveness of multidata causal discovery for hurricane intensity prediction
Figure 3: Example results for the 24 hour intensity change forecast (DELV24) from Fold 3 using the SHIPS+ERA5 predictor set for SHIPS predictors. (a) Coefficient of determination $R^{2}$ on training, validation, and test sets plotted against the number of selected predictors, each point corresponding to a different value of the M-PC causal discovery hyperparameter $pc_{\alpha}$ (bottom scale). The vertical dashed line indicates the configuration with the highest validation $R^{2}$. (b) Variable selection abacus: each dot shows the presence of a predictor across the $pc_{\alpha}$ range in the test set. Variables are colored by group (e.g., Original SHIPS predictors, Shear, Humidity), vertical dashed line marks the best validation score, and encircled dots highlight the occurrence of new shortlisted predictors for SHIPS.
We now present results from the ERA5-based replication experiments, focusing on how the M-PC framework operates and how the predictor selections translate into statistical forecast skill using MLR. Figure 3 illustrates one realization of the ERA5-based causal discovery experiments for the 24 hour intensity change forecast (DELV24), shown for a representative cross-validation fold (Fold 3) under the noASSUMPS (“kitchen-sink”) configuration. The top panel shows training, validation, and test $R^{2}$ scores as a function of the number of input predictors in the MLR, controlled by the M-PC hyperparameter ${pc}_{\alpha}$. Higher values of ${pc}_{\alpha}$ correspond to a less strict statistical threshold for rejecting independence, allowing more predictors to be included in the causal model whereas a lower ${pc}_{\alpha}$ is more strict, resulting in a more stringent selection of predictors.
@@
The bottom panel presents the variable selection abacus, highlighting predictor presence and groupings across the range of ${pc}_{\alpha}$ values. The encircled dots highlight the predictors included in the final shortlisted set, identified based on their consistent selection frequency across multiple folds. The same procedure was repeated for seven cross-validation folds, all forecast lead times, and both experimental setups. Given the volume of results, Figure 3 presents one representative fold for the noASSUMPS experiment, while the best performing folds for lead times of up to 120 hours are provided in the Supplementary Information (Figures S1–S10). The best potential causal predictors that complement or improve upon the baseline SHIPS predictors for TC intensity change forecasting tend to be consistently shortlisted by the causal feature selection framework across different data splits and validated configurations, ensuring that the predictors perform well over a wide range of TC cases.